# NeRF-To-Real Tester: Neural Radiance Fields as Test Image Generators for Vision of Autonomous Systems

### v1

@@ Abstract
Index Terms:
neural radiance fields, autonomous systems, metamorphic testing
@@ I Introduction
“I am not crazy; my reality is just different from yours.” says the Cheshire Cat [1]. A roboticist might conclude that the Cheshire cat experiences the simulation-to-reality gap; the gap between the confines of system’s own objective reality, and the complexities of the real world. For Autonomous Underwater Vehicles (AUVs) and Unmanned Aerial Vehicles (UAVs) overfitting on simulation-based testing setups fundamentally leads to poor performance in the operating conditions of the real world. Testing research aims to decrease the gap.
@@
Autonomous inspection of infrastructure both underwater and on land is a quickly growing market, with example applications including surveying buildings and bridges, processing plants, and tracking environmental changes in on- and off-shore construction sites or wind energy farms. Robotic inspections facilitate automatic detection of material degradation, initialisation of preventative measures to avoid equipment damage, as well as scheduling timely maintenance operations. To enable safe autonomous operations in real-world environments with potentially adverse weather conditions such as wind or underwater currents, we need to establish the quality of different algorithmic implementations used for navigation, and ensure their reliability of establishing a consistent understanding of their environment.
@@
AUV missions can extend over long time and distance, which increases the risk of drift and pose estimation errors due to underwater currents and sensor failures [2]. Integrating image-based perception into the navigation stack can reduce these failures and enhance the AUV’s real-time understanding of the scene. Consider an example subsystem in a typical vision stack: An Interest Point Detector (IPD) is used during visual simultaneous localisation and mapping (vSLAM) to track the AUV’s position in an open environment. Like many vision components, IPDs are negatively affected by the adverse conditions underwater such as bad visibility or sudden positional changes due to underwater currents. Similarly, in aerial surveillance, accurate object detection and classification are crucial for safe mission completion. Light-weight Deep Neural Networks (DNNs) can be deployed on modern UAVs, such as quadrocopters, to run real-time image classification, facilitating adaptive decision making for object tracking and obstacle avoidance. Sudden adverse wind turbulence can cause an UAV to fall short in their understanding of the scene. Unfortunately, despite the importance of IPDs and classifiers for safety of the autonomous vehicles, existing methods for testing such components fall short in generating real and diverse image data that exposes the system under test to realistic dynamics.
@@
Figure 1: Trade-offs between the realness and amount of data.
@@
For generating test data, simulators offer controlled environments and the ability to manipulate lighting conditions and scene characteristics. However, their effectiveness is constrained by the expertise and domain knowledge of the engineers who design them. A good performance on a simulation benchmark does not necessarily guarantee a good performance in the real world, hence the expression “simulation-to-reality gap”. Additionally, simulators often fail to encompass the full spectrum of unexpected safety-critical scenarios that may arise in real-world environments. Consequently, there is a pressing need for more diverse and realistic test data that accurately represents the challenges faced both by AUVs and UAVs. A trend in 3D scene reconstruction has emerged in recent years known as Neural Radiance Fields (NeRFs) [3], which establish a differentiable environmental representation by training a fully connected network on camera images and their poses. This allows to render new views from the learned scene, by querying a camera pose. This positions NeRFs as a source of more realistic test data when compared to manually engineered simulations, placing them at a sweet spot between simulated and real data when it comes to amount and cost of obtaining data (Fig. 1).
@@
  • An adaptation of metamorphic testing to testing vision components in 3D space on real and NeRF-generated data (Sect. IV),
@@ II Background
Neural Radial Fields [3] are function approximations of the form $\theta:(\mathbb{R}^{3},\mathbb{R}^{2})\rightarrow(\mathbb{R}^{3},\mathbb{R})$. Given a set of images with associated 3D locations $\mathbf{x}$ and viewing angles $\mathbf{d}$ of one particular scene or object, a NeRF learns to map $\theta(\mathbf{x},\mathbf{d})\rightarrow(\mathbf{c},\sigma)$, where the output $\mathbf{c}$ is the emitted RGB colour and $\sigma$ is the volume density. See Fig. 2 for an overview of the NeRF training and rendering process. A trained NeRF model is able to render new views of a scene from previously unseen viewpoints, i.e. combinations of ($\mathbf{x},\mathbf{d}$). This is achieved by standard volumetric rendering techniques, i.e. a ray $\mathbf{r}$ is “fired” from the camera origin $\mathbf{o}$, as shown in Fig. 2. When tracing the ray between the near and far bounds $t_{n}$, $t_{f}$ and taking samples along the ray at different points $t$, the expected colour $C(\mathbf{r})$ of this ray $\mathbf{r}(t)=\mathbf{o}+t\mathbf{d}$ starting at the camera origin $\mathbf{o}$ is given by:
@@
The function $\textrm{Tr}(t)$ represents the probability of the ray travelling through $\sigma$ without being obstructed by any particle:
@@
Note that NeRF models are unusual comparing to most other deep learning models; instead of learning patterns in a data set to enable generalisations or classifications the aim here is to learn a 3D scene representation by overfitting to a particular single scene or object.
@@
Figure 2: (Top) We take a set of real images of a scene $i_{\textrm{real}}\!\in\!S$ as input. The NeRF model $\theta$ learns to map the estimated camera poses $\mathbf{x},\mathbf{d}$ (yellow) to colour $\mathbf{c}$ and volumetric density $\sigma$. (Bottom) We pass new camera poses $(\mathbf{x},\mathbf{d})$ to the NeRF model $\theta$, creating previously unseen views $i_{\textrm{nerf}}$ of the scene. Camera rays $\mathbf{r}(t)$ (red) are projected from the camera origin along the trajectory between a near and far bounds $t_{n}$ and $t_{f}$.
@@ III What is a good test image?
Producing valid test images for perception subsystems is an active and challenging research area. It involves generating input images that preserve photometric, semantic, geometric and other contextual information while introducing perceivable, realistic as well as diverse changes, for which we are still able to determine the expected test output. Especially for navigation applications such as vSLAM, spatio-temporal consistency is vital to preserve geometric information across consecutive frames under pose translations. Therefore, we outline three key requirements for the generation of test images, that allow us to confidently assess the reliability of a system in real-world scenarios:
@@
These are interconnected requirements on test image data that ensure a robust and comprehensive testing of perception systems for navigation. Without realness we cannot confidently assess the reliability of a system in real-world scenarios. Without diversity we cannot test generalisation. Without spatio-temporal consistency, testing of a navigational systems is void.
@@
Strictly speaking, valid test inputs for testing perception are the images of real world scenes as captured by a representative camera. Methods like cross-validation enable model evaluation with images from operational conditions. However, using only real world scene data is often too costly. For instance, in underwater robotics, acquiring data necessitates offshore missions, involving considerable resources and time. Therefore, alternative methods are necessary to obtain additional yet realistic images. One possibility is to resort to simulation [4]. Unfortunately, the images produced by simulators are still not sufficiently realistic. They are presently more useful for pre-training perception models, and for visual assessment of robot actions by humans. Adversarial methods create imperceptibly small noise vectors that, when added to a real image, push it over a decision boundary [5]. Similarly, coverage-based methods that generate images triggering faulty behaviour, are inherently adversarial, focusing on the model rather than the operational domain distribution [6, 7, 8]. The adversarial inputs are realistic from the human eye perspective and may pose a significant threat from a security point of view. However, they are not realistically producible by representative cameras in operational scenes, so they are not interesting for testing robustness of perception. For these reasons, we study NeRFs, a model designed for realistic representation of scenes.
@@
Test data generation aims to extend beyond the training data, while maintaining realness. Simple methods like rotating and scaling images, known as data augmentation, are common in DNN training. Most of the existing research on the diversity of the generated data focuses on more advanced transformation techniques and on measuring realness or validity of the generated inputs. One group of methods patches existing images; pasting new objects and changing properties of objects (e.g. lighting or colour [9]). However, these can be easily identified as unreal by humans. A less direct approach is to perform search in the feature space [10, 11], or to train Generative Adversarial Networks (GANs) which produce images that appear realistic. GANs have also been used to change global properties of an image, like weather conditions [12]. A GAN incorporates a discriminator network which optimises the realness of the generated images. Crucially, the discriminator network learns statistical patterns and not the semantic content in the generated images. It might learn to exploit weaknesses in the discriminator without actually learning how to represent the underlying distribution and therefore alternative methods are still needed. Other methods such as training Variational Autoencoders (VAEs) [13] and using their reconstruction probabilities as proof for validity is not effective because each model’s output is conditioned on its own dataset’s characteristics. Comparing these probabilities are not directly comparable since they are relative to their underlying distributions. Proxy metrics are typically used to evaluate realness of images produced by generative models. Peak-Signal-To-Noise Ratio (PSNR) is based on the mean squared error, and therefore sensitive to rotations, which otherwise have little impact on human perception of realness. Structural Similarity Index (SSIM) estimates structural similarity between two images according to luminance and contrast [14]. Learned Perceptual Image Patch Similarity (LPIPS) measures similarity in a DNN feature representation [15]. Unlike the above, we use transformations in the NeRF space (rolls and translations) to create new inputs for the perception components. Instead of challenging the SUT with new object and lighting configurations, we generate new view angles and new occlusions.
@@
For testing perception in the context of navigation, one needs to produce pairs, or sequences, of images that differ in the camera location, and that manifest the corresponding changes to the geometry of objects in the scene—consistent variation of positions, view angles, and occlusions. NeRFs were originally developed precisely for this requirement. Unlike GANs, where the learned distribution is a complex manifold in a high-dimensional space, giving poor control in the image generation process, NeRFs can render new images parameterised with camera positions in Euclidian space. This allows to emulate the motion of the agent, while maintaining the realness driven by the training data in the scene.
@@ IV Method
We describe our method to use NeRFs as test image generators within a customised metamorphic testing procedure.
Figure 3: Metamorphic testing: We establish metamorphic relations (MR) between pairs of test inputs and between the corresponding outputs. The test fails if the latter do not hold.
@@ IV-A Metamorphic Testing
Metamorphic Testing (MT) is a testing framework to assess the accuracy and robustness of a System Under Test (SUT) [16]. It is particularly well-suited when the quality of the expected output of a SUT is difficult to establish due to high dimensionality of the test data or absence of ground truth. Unlike traditional testing methods that rely on input-output pairs, in MT we formulate metamorphic relations (MRs) between input-input and output-output pairs that the SUT is expected to comply with (Fig. 3). MT therefore aims to capture the intrinsic properties of the data that the system processes. While other testing techniques aim at detecting bugs on a code level, MT can uncover critical inconsistencies on a behavioural level of a SUT by observing discrepancies in test outputs, given interrelated test inputs. MT is highly suitable for testing image-based perception systems where the inherent complexity and unstructured nature of image data poses significant challenges. Given a SUT $f$ that processes images $i$ of a scene $S$, a metamorphic relation can be expressed as:
$\forall i\in S.\;f(g_{I}(i))=g_{O}(f(i))\enspace.$ (3)
Here, $g_{I},g_{O}$ are functions that transform the programme input and output domain of $f$ respectively. We transform the test images $i$ according to function $g_{I}$ to generate new tests and validate them against the output of the SUT $f(i)$ by function $g_{O}$. In this work, we utilise NeRFs as automatic test image generators $g_{I}$ while simultaneously monitoring the effects that NeRF-generated images have on the output of a SUT (Fig. 4). To achieve this we employ a range of transformations of an input image as our input MR and evaluate discrepancies between the outputs of the SUT for the transformed images using performance metrics (Fig. 5). The strength of NeRFs lies in their ability to render realistic images from previously unseen camera poses $(\mathbf{x},\mathbf{d})$ of a scene or object, while preserving its photometric, structural and geometric information, which simultaneously offers image diversity by giving different viewpoints. We exploit this ability by applying a range of transformations $T$ to $(\mathbf{x},\mathbf{d})$ in the 3D NeRF space. We denote the a transformed image $i_{\tau}$. We report how a SUT $f$ responds to a range of diverse test images by the amount of $\delta(f(i_{\textrm{nerf}}),f(i_{\tau}))$. The predicate INC below captures the violation written using a performance metric $q$ into real numbers:
$\textrm{INC}(f)=q(f(i_{\textrm{nerf}}),f(i_{\tau}))>\epsilon$ (4)
Figure 4: N2R-Tester: (1) an image processing SUT $f$, (2) a metamorphic relation (MR) in form of a pose transformation $\tau$ rendered in a NeRF model, (3) an MR $\delta$ checking inconsistencies between the outputs $f(i_{\textrm{nerf}})$, $f(i_{\tau})$ of the SUT.
@@ IV-B NeRF-To-Real Tester
N2R-Tester $\langle f,S,T,\delta\rangle$ comprises a system under test $f$, a set of images $S$ of a scene or object, a list of pose transformations $T$, and metamorphic relations $\delta$. N2R-Tester takes the image set $S$ and reconstructs a 3D scene solely based on these 2D input images in form of a NeRF model. We achieve this by feeding all images $i\in S$ to a structure-from-motion (SfM) component, colmap [17], which selects a subset of all images $S_{\textrm{real}}\subseteq S$ to get estimated camera poses $(\mathbf{x},\mathbf{d})$ for each image in $i_{\textrm{real}}\in S_{\textrm{real}}$. These image/camera-pose combinations are used to train a NeRF model $\theta$. Now we render the same camera poses in the NeRF model to get all corresponding NeRF-generated images $i_{\textrm{nerf}}$ from the identical viewpoints. Subsequently, we apply a number of different transformations $\tau\in T$ to each of the camera poses $(\mathbf{x},\mathbf{d})$ to render the transformed images in the NeRF model to get $i_{\tau}$. The supported transformations are translations of position as well as rotations in roll, pitch and yaw angle.
Figure 5: N2R-Tester Process: Given a set of real images of a scene $S_{\textrm{real}}$ we train a NeRF model $\theta$. By designing a set transformations $T$ we can generate a high number of additional test images and check the consistency of the SUT behaviour.

### v2

@@ I Introduction
“I am not crazy; my reality is just different from yours.” says the Cheshire Cat [1]. A roboticist might conclude that the Cheshire cat experiences the simulation-to-reality gap; the gap between the confines of system’s own objective reality, and the complexities of the real world. For Autonomous Underwater Vehicles (AUVs) and Unmanned Aerial Vehicles (UAVs), overfitting to simulation-based testing setups can lead to poor performance in the operating conditions of the real world. Testing research aims to decrease this gap.
@@
Autonomous inspection of infrastructure, both underwater and on land, is a fast growing market. Applications include surveying constructions sites or wind farms for facilitating automatic detection of material degradation or environmental changes. To enable safe autonomous operations in real-world environments and sudden positional jumps due to adverse weather conditions such as wind or underwater currents, we need to establish the quality of different navigation systems and implementations, and ensure their reliability of establishing a consistent understanding of their environment.
@@
Integrating image-based perception into an autonomous navigation stack can enhance the AUV/UAV’s real-time understanding of a scene for more fine-grained tasks. Light-weight Deep Neural Networks (DNNs) are often the default choice for tasks such as object detection or visual simultaneous localization and mapping (vSLAM). In aerial surveillance, accurate object classification and tracking are crucial for facilitating adaptive decision making. In underwater inspections, an Interest Point Detector (IPD) serves as the initial system to track the AUV’s position in an open environment. Both image classifiers and IPDs are negatively affected by the adverse conditions such as bad visibility or sudden positional changes due to underwater currents or wind turbulence in UAV and AUVs respectively.
@@
Figure 1: Trade-offs between the data realness and amount/cost.
@@
For generating test data, simulators offer controlled environments and the ability to manipulate lighting conditions and scene characteristics. However, their effectiveness is constrained by the expertise and domain knowledge of the engineers who design them. A good performance on a simulation benchmark does not necessarily guarantee a good performance in the real world. Consequently, there is a pressing need for more diverse and realistic test data that accurately represents the challenges faced both by AUVs and UAVs. A trend in 3D scene reconstruction has emerged in recent years known as Neural Radiance Fields (NeRFs) [2], providing a 3D representation of a scene in form of a neural network trained on camera poses and their respective images. NeRFs allow rendering novel views from the learned scene, by querying previously unseen camera poses. This positions NeRFs as a source of more realistic test data compared to engineered simulations, placing them at a sweet spot considering the trade-off when it comes to amount and cost of obtaining data (Fig. 1).
@@
  • An adaptation of metamorphic testing for vision components in 3D space on real and NeRF-generated data (Sect. IV),
@@
To the best of our knowledge, we are the first to investigate how NeRFs can be applied in test image generation.
@@ II Background
Neural Radial Fields [2] (NeRFs) are function approximations of the form $\theta:(\mathbb{R}^{3},\mathbb{R}^{2})\rightarrow(\mathbb{R}^{3},\mathbb{R})$. A NeRF learns a scene by mapping 3D camera locations $\mathbf{x}$ and viewing angles $\mathbf{d}$ to RGB color $\mathbf{c}$ and volume density $\sigma$, thus $\theta(\mathbf{x},\mathbf{d})\rightarrow(\mathbf{c},\sigma)$. A trained NeRF model is able to render novel views of a scene from previously unseen viewpoints, i.e. combinations of ($\mathbf{x},\mathbf{d}$), as seen in Fig. 2. This is achieved by standard volumetric rendering techniques, i.e. a ray $\mathbf{r}$ is “fired” from the camera origin $\mathbf{o}$, as shown in Fig. 2. When tracing the ray between the near and far bounds $t_{n}$, $t_{f}$ and taking samples along the ray at different points $t$, the expected color $C(\mathbf{r})$ of this ray $\mathbf{r}(t)=\mathbf{o}+t\mathbf{d}$ starting at the camera origin $\mathbf{o}$ is given by:
@@
The function $\textrm{Tr}(t)$ represents the probability of the ray traveling through $\sigma$ without being obstructed by any particle:
@@
Figure 2: (Top) We take a set of real images of a scene $i_{\textrm{real}}\!\in\!S$ as input. The NeRF model $\theta$ learns to map the estimated camera poses $\mathbf{x},\mathbf{d}$ (yellow) to colour $\mathbf{c}$ and volumetric density $\sigma$. (Bottom) We pass new camera poses $(\mathbf{x},\mathbf{d})$ to the NeRF model $\theta$, creating previously unseen views $i_{\textrm{nerf}}$ of the scene. Camera rays $\mathbf{r}(t)$ (red) are projected from the camera origin along the trajectory between a near and far bounds $t_{n}$, $t_{f}$.
@@ III What is a good test image?
For autonomous systems that operate in the real world, producing valid test images for perception is an active and challenging research area. It involves generating input images that preserve photometric, semantic, geometric and other contextual information while introducing perceivable, realistic as well as diverse changes, for which we are still able to determine the expected test output. Especially for navigation applications such as vSLAM, spatio-temporal consistency is vital to preserve geometric information across consecutive frames under pose translations. We outline three key requirements for the generation of test images, that allow us to confidently assess the reliability of a system in real-world scenarios:
@@
These are interconnected requirements on test image data that ensure a robust and comprehensive testing of perception systems for navigation. Without realness we cannot confidently assess the reliability of a system in real-world scenarios. Without diversity we cannot test generalization. Without spatio-temporal consistency, testing of a navigational systems is void.
@@
Strictly speaking, valid test inputs for testing perception are the images of real world scenes captured by a camera. However, using only real world scene data is often too costly. For instance, in underwater robotics, acquiring data necessitates offshore missions, involving considerable resources and time. Therefore, alternative methods are necessary to obtain additional yet realistic images. One possibility is to resort to simulation [3]. The images produced by simulators often possess an artificial visual quality and intensify the simulation-to-reality gap. Alternatively, adversarial methods create imperceptibly small noise vectors that, when added to a real image, push it over a decision boundary [4]. Similarly, coverage-based methods that generate images triggering faulty behavior, are inherently adversarial, directly attacking a model based on its internal learned representations [5, 6]. We treat the perception systems as black-box and therefore provide a more generalizable method.
@@
Test data generation aims to extend beyond the training data, while maintaining realness. Simple methods like rotating and scaling images, known as data augmentation, are common in DNN training. Most of the existing research on the diversity of the generated data focuses on more advanced transformation techniques and on measuring realness or validity of the generated inputs. One group of methods patches existing images; pasting new objects and changing properties of objects (e.g. lighting or color [7]). However, these can be easily identified as unreal by humans. A less direct approach is to perform search in the feature space [8], or to change global properties of an image, like weather conditions, by use of Generative Adversarial Networks (GANs) [9]. A GAN incorporates a discriminator network which optimizes the realness of the generated images. Crucially, the discriminator network learns statistical patterns and not the semantic content in the generated images. It might learn to exploit weaknesses in the discriminator without actually learning the representation of the underlying distribution. Unlike the above, we use transformations in the NeRF space (rolls and translations) to create new inputs for the perception components. Instead of challenging the SUT with new object and lighting configurations, we generate new view angles and new occlusions.
@@
For testing perception in the context of navigation, one needs to produce image sequences that differ in the camera location, and that manifest the corresponding changes to the geometry of objects in the scene-consistent variation of positions, view angles, and occlusions. NeRFs were originally developed precisely for this requirement. Unlike GANs, where the learned distribution is a complex manifold in a high-dimensional space, giving poor control in the image generation process, NeRFs can render new images parameterized with camera positions in Euclidian space. This allows to emulate the motion of the agent, while maintaining the realness driven by the training data in the scene.
@@ IV Method
We describe our method to use NeRFs as test image generators within a customized metamorphic testing procedure.
Figure 3: Metamorphic testing: Metamorphic relations (MRs) are established between pairs of test inputs and between the corresponding outputs. The test fails if MR2 does not hold.
@@ IV-A Metamorphic Testing
Metamorphic Testing (MT) is a testing framework to assess the accuracy and robustness of a System Under Test (SUT) [10]. It is particularly well-suited when the quality of the expected output of a SUT is difficult to establish due to high dimensionality of the test data or absence of ground truth. Unlike traditional testing methods that rely on input-output pairs, in MT we formulate metamorphic relations (MRs) between input-input and output-output pairs that the SUT is expected to comply with (Fig. 3). MT therefore aims to capture the intrinsic properties of the data that the system processes. While other testing techniques aim at detecting bugs on a code level, MT can uncover critical inconsistencies on a behavioral level of a SUT by observing discrepancies in test outputs, given interrelated test inputs. MT is highly suitable for testing image-based perception systems where the inherent complexity and unstructured nature of image data poses significant challenges. Given a SUT $f$ that processes images $i$ of a scene $S$, a metamorphic relation can be expressed as $\forall i\in S.\;f(g_{I}(i))=g_{O}(f(i))$. Here, $g_{I},g_{O}$ are functions that transform the program input and output domain of $f$ respectively. We transform the test images $i$ according to function $g_{I}$ to generate new tests and validate them against the output of the SUT $f(i)$ by function $g_{O}$. In this work, we utilize NeRFs as automatic test image generators $g_{I}$ while simultaneously monitoring the effects that NeRF-generated images have on the output of a SUT (Fig. 4). To achieve this we employ a range of transformations of an input image as our input MR and evaluate discrepancies between the outputs of the SUT for the transformed images using critical performance metrics. The strength of NeRFs lies in their ability to render realistic images from previously unseen camera poses $(\mathbf{x},\mathbf{d})$ of a scene or object, while preserving its photometric, structural and geometric information, which simultaneously offers image diversity by giving different viewpoints. We exploit this ability by applying a range of transformations $T$ to $(\mathbf{x},\mathbf{d})$ in the 3D NeRF space. We denote the a transformed image $i_{\tau}$. We report how a SUT $f$ responds to a range of diverse test images by the changes in MR outputs, i.e. $\delta(f(i_{\textrm{nerf}}),f(i_{\tau}))$. The predicate INC captures the violation written using a performance metric $q$ into real numbers:
$\textrm{INC}(f)=q(f(i_{\textrm{nerf}}),f(i_{\tau}))>\epsilon.$ (3)
Figure 4: N2R-Tester: (1) an image processing SUT $f$, (2) a metamorphic relation (MR) in form of a pose transformation $\tau$ rendered in a NeRF model, (3) an MR $\delta$ checking inconsistencies between the SUT outputs $f(i_{\textrm{nerf}})$ and $f(i_{\tau})$.
@@ IV-B NeRF-To-Real Tester
N2R-Tester $\langle f,S,T,\delta\rangle$ comprises a system under test $f$, a set of images $S$ of a scene or object, a list of pose transformations $T$, and metamorphic relations $\delta$. N2R-Tester takes the image set $S$ and reconstructs a 3D scene solely based on these 2D input images in form of a NeRF model. We achieve this by feeding all images $i\in S$ to a structure-from-motion tool, colmap [11], which selects a subset of all images $S_{\textrm{real}}\subseteq S$ to get estimated camera poses $(\mathbf{x},\mathbf{d})$ for each image in $i_{\textrm{real}}\in S_{\textrm{real}}$. These image/camera-pose combinations are used to train a NeRF model $\theta$. Now we render the same camera poses in the NeRF model to get all corresponding NeRF-generated images $i_{\textrm{nerf}}$ from the identical viewpoints. Subsequently, we apply a number of different transformations $\tau\in T$ to each of the camera poses $(\mathbf{x},\mathbf{d})$ to render the transformed images in the NeRF model to get $i_{\tau}$. The supported transformations are translations of position as well as rotations in roll, pitch and yaw angle.